# Non-Radial Solution Structures for Quasilinear Hamilton–Jacobi–Bellman Equations in Bounded Settings

### v1

@@ 1 Introduction
where $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded $C^{2}$ convex domain with smooth boundary $\partial\Omega$, $\sigma>0$ represents the diffusion coefficient, $\alpha\in(1,2]$ is the cost exponent, and
@@
The motivation for considering (1.1) is twofold. First, the equation arises in image restoration problems, as discussed in our recent work [2], where the parameter $\alpha$ is used for contrast of a image (as a constraint) in the image model. Second, from a purely mathematical perspective, the case $\alpha\in(1,2)$ remains largely unexplored for general bounded domains $\Omega$ (see [2, 3]).
@@
Let $h:\overline{\Omega}\rightarrow[0,\infty)$ be continuous and sub-quadratic (i.e., there exist constants $C_{h}>0$ and $q\in[0,2)$ such that $h(y)\leq C_{h}(1+|y|^{q})$ for all $y\in\overline{\Omega}$), and let $g\in C(\overline{\Omega},\left[0,\infty\right))$ with $g\geq 0$ be the boundary datum. Then the Dirichlet problem (1.1) admits a unique positive solution $V\in C^{2}(\Omega)\cap C(\overline{\Omega})$. Moreover, if $g>0$ on $\partial\Omega$ and $h>0$ in $\Omega$, then $V>0$ in $\overline{\Omega}$.
@@
The sub-quadratic growth condition on $h$ is optimal for this class of problems. For $q\geq 2$, the gradient term may dominate and solutions may fail to exist globally. The regularity assumption $h\in C(\overline{\Omega})$ can be relaxed to $h\in L^{\infty}(\Omega)$ with appropriate modifications to the proof.
@@ 2.1 Step 1: Construction of ordered positive sub- and super-solutions
$c_{1}:=\min_{x\in\overline{\Omega}}g(x),\qquad V_{-}(y):=c_{1},\quad y\in\overline{\Omega}.$
@@
because $h\geq 0$. Hence $V_{-}$ is a (classical) sub-solution in $\Omega$, and $V_{-}\leq g$ on $\partial\Omega$. Since $c_{1}\geq 0$, $V_{-}$ is positive.
Super-solution. Let
$c_{2}:=\max_{x\in\overline{\Omega}}g(x),$
fix a constant $B>0$ to be determined, and define
$V_{+}(y):=c_{2}+B\,\phi(y),\qquad y\in\overline{\Omega}.$ (2.2)
By construction, $V_{+}=c_{2}\geq g$ on $\partial\Omega$ since $\phi=0$ there, and $V_{+}\geq c_{2}\geq g\geq 0$ in $\Omega$ because $\phi>0$ in $\Omega$. Computing the derivatives:
@@
Thus $V_{+}$ is a classical super-solution to (1.1), satisfies $V_{+}\geq g$ on $\partial\Omega$, and $V_{+}>0$ in $\Omega$ (since $\phi>0$ in $\Omega$ and $c_{2}\geq 0$). Order and positivity. By construction,
$V_{-}(y)=c_{1}\;\leq\;c_{2}+B\,\phi(y)=V_{+}(y)\quad\text{for all }y\in\Omega,$
with equality on $\partial\Omega$ since $\phi=0$ there and $c_{1}\leq g(x)\leq c_{2}$ for $x\in\partial\Omega$. Both $V_{-}$ and $V_{+}$ are positive provided $g\geq 0$.
@@
$V_{-}(y)\equiv c_{1}\quad\text{and}\quad V_{+}(y)=c_{2}+B\phi(y)\quad\text{for }B\geq 2H/\sigma^{2},$
@@
  • If one prefers a radial barrier and $\Omega$ is $C^{2}$ and strictly contained in a ball $B_{R}(0)$, one may also take $V_{+}(y)=c_{2}+B\psi(y)$ with $\psi$ solving $\Delta\psi=1$ in $\Omega$, $\psi=0$ on $\partial\Omega$ (as above), which avoids geometric mismatch of a purely radial function with a general $\partial\Omega$.
@@ 2.4 Step 4: Uniqueness
    Initialize $V^{(0)}=c_{2}+B\phi$ with $B\geq 2H/\sigma^{2}$.
@@ 3.1.2 Exit time and cost functional
Fix a running cost $h:\overline{\Omega}\rightarrow[0,\infty)$ that is continuous and sub-quadratic, and a boundary cost $g:\overline{\Omega}\rightarrow[0,\infty)$ that is continuous. For each $x\in\Omega$ and admissible control $v\in\mathcal{U}_{\text{ad}}$, define the cost functional
$J(x;v)\;:=\;\mathbb{E}\!\left[\int_{0}^{\tau^{x,v}}\Big(h(X_{t}^{x,v})+|v_{t}|^{\alpha}\Big)\,dt+g(X_{\tau^{x,v}}^{x,v})\right].$ (3.5)
The terminal cost $g(X_{\tau}^{x,v})$ represents the cost incurred upon exiting the domain.
@@ 3.1.3 Dynamic programming and verification
Let $V\in C^{2}(\Omega)\cap C(\overline{\Omega})$ solve the HJB equation (3.14) with boundary condition $V=g$ on $\partial\Omega$, and assume that $V$ satisfies polynomial growth. Define the feedback control $v^{*}(x)$ by (3.17). Then $V(x)=J(x;v^{*})$ is the value function, and $v^{*}$ is optimal.
@@
Under the feedback control $v^{*}$, the process $M_{t}$ defined in (3.7) becomes a local martingale. By the polynomial growth of $V$ and the sub-quadratic growth of $h$, $M_{t}$ is a true martingale. At the exit time $\tau$, we have $V(X_{\tau})=g(X_{\tau})$ by the boundary condition. Taking expectations in $M_{0}=\mathbb{E}[M_{\tau}]$ yields
$V(x)=\mathbb{E}\!\left[\int_{0}^{\tau}\big(|v_{t}^{\ast}|^{\alpha}+h(X_{t}^{x,v^{*}})\big)\,dt+g(X_{\tau}^{x,v^{*}})\right]=J(x;v^{\ast}).$
For any other admissible control $v$, the supermartingale property gives $V(x)\leq J(x;v)$, confirming optimality of $v^{*}$. ∎
@@ 5.2 Software implementation
For the special case of radially symmetric problems ($\Omega=B_{R}(0)$ and $h$ radial), numerical codes implementing the above algorithm are available in our companion repository [2]. The codes solve the radial ODE (4.1) using finite difference methods with adaptive mesh refinement.

### v2

# Radial and Non-Radial Solution Structures for Quasilinear Hamilton–Jacobi–Bellman Equations in Bounded Settings

@@ 1 Introduction
where $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded $C^{2}$ convex domain with smooth boundary $\partial\Omega$, $\sigma>0$ represents the diffusion coefficient, $\alpha\in(1,2]$ is the cost exponent, and $g\geq 0$ is a constant,
@@
The motivation for considering (1.1) is twofold. First, the equation arises in image restoration problems, as discussed in our recent work [2], where the parameter $\alpha$ close to $1$ is used for contrast of a image (as a constraint) in the image model. Second, from a purely mathematical perspective, the case $\alpha\in(1,2)$ remains largely unexplored for general bounded domains $\Omega$ (see [2, 3]).
@@
Let $h:\overline{\Omega}\rightarrow[0,\infty)$ be continuous and sub-quadratic (i.e., there exist constants $C_{h}>0$ and $p:=\frac{\alpha}{\alpha-1}\in[2,\infty)$ with $\alpha\in(1,2]$ such that $h(y)\leq C_{h}(1+|y|^{p})$ for all $y\in\overline{\Omega}$), and let $g\in[0,\infty)$ be a constant boundary datum. Then the Dirichlet problem (1.1) admits a unique positive solution $V\in C^{2}(\Omega)\cap C(\overline{\Omega})$. Moreover, if $g>0$ on $\partial\Omega$ and $h>0$ in $\Omega$, then $V>0$ in $\overline{\Omega}$.
@@
The regularity assumption $h\in C(\overline{\Omega})$ can be relaxed to $h\in L^{\infty}(\Omega)$ with appropriate modifications to the proof.
@@ 2.1 Step 1: Construction of ordered positive sub- and super-solutions
$V_{-}(y):=g,\quad y\in\overline{\Omega}.$
@@
because $h\geq 0$. Hence $V_{-}$ is a (classical) sub-solution in $\Omega$, and $V_{-}=g$ on $\partial\Omega$. Since $g\geq 0$, $V_{-}$ is positive.
Super-solution. Fix a constant $B>0$ to be determined, and define
$V_{+}(y):=g+B\,\phi(y),\qquad y\in\overline{\Omega}.$ (2.2)
By construction, $V_{+}=g$ on $\partial\Omega$ since $\phi=0$ there, and $V_{+}\geq g\geq 0$ in $\Omega$ because $\phi>0$ in $\Omega$. Computing the derivatives:
@@
Thus $V_{+}$ is a classical super-solution to (1.1), satisfies $V_{+}=g$ on $\partial\Omega$, and $V_{+}>0$ in $\Omega$ (since $\phi>0$ in $\Omega$ and $g\geq 0$).
Order and positivity. By construction,
$V_{-}(y)=g\;\leq\;g+B\,\phi(y)=V_{+}(y)\quad\text{for all }y\in\Omega,$
with equality on $\partial\Omega$ since $\phi=0$ there. Both $V_{-}$ and $V_{+}$ are positive provided $g\geq 0$.
@@
$V_{-}(y)\equiv g\quad\text{and}\quad V_{+}(y)=g+B\phi(y)\quad\text{for }B\geq 2H/\sigma^{2},$
@@
  • If one prefers a radial barrier and $\Omega$ is $C^{2}$ and strictly contained in a ball $B_{R}(0)$, one may also take $V_{+}(y)=g+B\psi(y)$ with $\psi$ solving $\Delta\psi=1$ in $\Omega$, $\psi=0$ on $\partial\Omega$ (as above), which avoids geometric mismatch of a purely radial function with a general $\partial\Omega$.
@@ 2.4 Step 4: Uniqueness
    Initialize $V^{(0)}=g+B\phi$ with $B\geq 2H/\sigma^{2}$.
@@ 3.1.2 Exit time and cost functional
Fix a running cost $h:\overline{\Omega}\rightarrow[0,\infty)$ that is continuous and sub-quadratic, and a constant boundary cost $g\in[0,\infty)$. For each $x\in\Omega$ and admissible control $v\in\mathcal{U}_{\text{ad}}$, define the cost functional
$J(x;v)\;:=\;\mathbb{E}\!\left[\int_{0}^{\tau^{x,v}}\Big(h(X_{t}^{x,v})+|v_{t}|^{\alpha}\Big)\,dt+g\right].$ (3.5)
The terminal cost $g$ represents the cost incurred upon exiting the domain.
@@ 3.1.3 Dynamic programming and verification
Let $V\in C^{2}(\Omega)\cap C(\overline{\Omega})$ solve the HJB equation (3.14) with constant boundary condition $V=g$ on $\partial\Omega$, and assume that $V$ satisfies polynomial growth. Define the feedback control $v^{\ast}(x)$ by (3.17). Then $V(x)=J(x;v^{\ast})$ is the value function, and $v^{\ast}$ is optimal.
@@
Under the feedback control $v^{\ast}$, the process $M_{t}$ defined in (3.7) becomes a local martingale. By the polynomial growth of $V$ and the sub-quadratic growth of $h$, $M_{t}$ is a true martingale. At the exit time $\tau$, we have $V(X_{\tau})=g$ by the boundary condition. Taking expectations in $M_{0}=\mathbb{E}[M_{\tau}]$ yields
$V(x)=\mathbb{E}\!\left[\int_{0}^{\tau}\big(|v_{t}^{\ast}|^{\alpha}+h(X_{t}^{x,v^{\ast}})\big)\,dt+g\right]=J(x;v^{\ast}).$
For any other admissible control $v$, the supermartingale property gives $V(x)\leq J(x;v)$, confirming optimality of $v^{\ast}$. ∎
@@ 5.2 Software implementation
For the special case of radially symmetric problems ($\Omega=B_{R}(0)$ and $h$ radial), numerical codes implementing the above algorithm are available in our companion repository [2]. See
https://github.com/coveidragos/code_contrast_image/blob/main/contrast_image.py
for the denoised-restored contrast scheme, or
https://github.com/coveidragos/code_contrast_image/blob/main/contrast_image_rest_denoised.py
for the restored-denoised contrast scheme.
The codes solve the radial ODE (4.1) using finite difference methods with adaptive mesh refinement.